2. Inputs and outputs Amys Performance Pizza is a small restaurant in Dallas that sells gluten-free pizzas. Amys very tinyProduction Function QUANTITY OF OUTPUT (Piezas) AAAA o 5 LABOR (Number of workers) Suppose that labor is Amys only variableUse the orange points (square symbol) to plot Amys total cost curve on the following graph using the quantities from the pre
Answer #1

In the short run, the workers are variable inputs and the ovens are fixed inputs.

A variable input is a resource or factor of production which can be changed in the short run by a firm as it seeks to change the quantity of output produced. A fixed input is a resource or factor of production which cannot be changed in the short run by a firm as it seeks to change the quantity of output produced.

The marginal product of labor (MPL) is the change in output that results from employing an added unit of labor.

MPL can be calculated as change in total output/ change in labor units.

Labor Output MPL Total Cost
0 0 0 20+0 =20
1 60 60 20+30(1)= 50
2 100 40 20+30(2)= 80
3 130 30 20+30(3)=110
4 150 20 20+30(4)= 140
5 160 10 20+30(5)= 170

Production function gives the technological relation between quantities of physical inputs and quantities of output of goods. It can derived by plotting output at various levels of labor units.

Total Cost= Fixed Cost + Variable Cost

Here, fixed cost = $20 and variable cost= $30 per worker

Thus Total cost= 20 + 30q where q represents units of labor

2005 PRODUCTION FUNCTION TOTAL 1407 120 에 1 2 3 4 5 Labor - 20 40 60 80 100 120 140 160 180

The law of diminshing marginal product of labor implies: (a) total output decreases at decreasing rate as you increase the quantity of labor.